# Localization and interpolation of parabolic $L^{p}$ Neumann problems

### v1

@@ 3. Basic estimates
The proof follows the strategy of [13, Lemma 4.1], adapted to the parabolic setting. Let $Q=Q_{r}(x,t)$ be a parabolic cube centered at the boundary. Without loss of generality, we can assume that the center $(x,t)=(0,0)\in\partial\Omega$. We assume that $u\in W^{1,2}(KQ\cap\Omega)$ is a weak solution to $Lu=0$ with zero Neumann data on $KQ\cap\partial\Omega$ for some $K\geq 1$ large (independent of $Q$ and $u$) to be determined during the proof, and we want to prove
(4.17) $\|\widetilde{N}(|\nabla u|{\mathds{1}}_{Q})\|_{L^{p}(\partial\Omega)}\lesssim r^{(n+1)/p}\Big(\mathchoice{{\vbox{\hbox{$\textstyle\mkern-1.5mu{-}\mkern-7.5mu{-}$}}\kern-4.675pt}}{{\vbox{\hbox{$\scriptstyle\mkern-1.5mu{-}\mkern-7.5mu{-}$}}\kern-4.02336pt}}{{\vbox{\hbox{$\scriptscriptstyle\mkern-1.5mu{-}\mkern-7.5mu{-}$}}\kern-2.53589pt}}{{\vbox{\hbox{$\scriptscriptstyle\mkern-1.5mu{-}\mkern-7.5mu{-}$}}\kern-1.7709pt}}\!\int\!\!\!\!\!\int_{KQ\cap\Omega}|\nabla u|^{2}dXdt\Big)^{\frac{1}{2}}.$
It is not hard to see that (4.16) follows from (4.17) by a covering argument.
We construct then $\varphi\in C^{\infty}_{0}(\mathbb{R}^{n+1})$ such that $\varphi\equiv 1$ on $\frac{3}{2}Q$, $\varphi\equiv 0$ outside $2Q$, $|\nabla\varphi|\lesssim\frac{1}{r}$, and $|\partial_{t}\varphi|\lesssim\frac{1}{r^{2}}$. By duality, i.e. (2.14), there exists $G\in L^{\infty}_{c}(\Omega,\mathbb{R}^{n})$ such that
@@
(4.19) $\|\widetilde{N}(\nabla u{\mathds{1}}_{Q})\|_{L^{p}(\partial\Omega)}\leq C\iint_{\Omega}{\mathds{1}}_{Q}\nabla u\cdot G\,dXdt$
Without loss of generality, we can assume that $G=G{\mathds{1}}_{Q}$. Let $c_{u}$ be a constant to be determined later, we have that
@@
$\|\widetilde{N}(\nabla u{\mathds{1}}_{Q})\|_{L^{p}(\partial\Omega)}\lesssim\iint_{\Omega}\nabla u\cdot G\,\varphi\,dXdt=\iint_{\Omega}\nabla(u-c_{u})\cdot G\,\varphi\,dXdt\\ =-\iint_{\Omega}(u-c_{u})\varphi\operatorname{div}G\,dXdt-\iint_{\Omega}(u-c_{u})\nabla\varphi\cdot G\,dXdt=:I_{0}+I_{1}$
The term $I_{1}$ is easy to treat. We choose $c_{u}=\mathchoice{{\vbox{\hbox{$\textstyle\mkern-1.5mu{-}\mkern-7.5mu{-}$}}\kern-4.675pt}}{{\vbox{\hbox{$\scriptstyle\mkern-1.5mu{-}\mkern-7.5mu{-}$}}\kern-4.02336pt}}{{\vbox{\hbox{$\scriptscriptstyle\mkern-1.5mu{-}\mkern-7.5mu{-}$}}\kern-2.53589pt}}{{\vbox{\hbox{$\scriptscriptstyle\mkern-1.5mu{-}\mkern-7.5mu{-}$}}\kern-1.7709pt}}\!\int\!\!\!\!\!\int_{4Q\cap\Omega}u\,dXdt$. By the bound $|\nabla\varphi|\lesssim r^{-1}$, we have that
$|I_{1}|\lesssim r^{n+1}\left(\mathchoice{{\vbox{\hbox{$\textstyle\mkern-1.5mu{-}\mkern-7.5mu{-}$}}\kern-4.675pt}}{{\vbox{\hbox{$\scriptstyle\mkern-1.5mu{-}\mkern-7.5mu{-}$}}\kern-4.02336pt}}{{\vbox{\hbox{$\scriptscriptstyle\mkern-1.5mu{-}\mkern-7.5mu{-}$}}\kern-2.53589pt}}{{\vbox{\hbox{$\scriptscriptstyle\mkern-1.5mu{-}\mkern-7.5mu{-}$}}\kern-1.7709pt}}\!\int\!\!\!\!\!\int_{2Q\cap\Omega}|G(X,t)|dXdt\right)\sup_{2Q\cap\Omega}|u-c_{u}|\lesssim r^{n+1}\left(\inf_{(x,t)\in 2Q\cap\partial\Omega}\widetilde{C}(\delta G)(x,t)\right)\sup_{2Q\cap\Omega}|u-c_{u}|$
by the definition of $\widetilde{\mathcal{C}}$. Since $u-c_{u}$ is also a solution to $Lu=0$ in $\Omega\cap KQ$ with zero Neumann data on $\partial\Omega\cap KQ$, we have by Lemma 3.3 and Lemma 3.5 that
(4.20) $\sup_{2Q\cap\Omega}|u-c_{u}|\leq C\mathchoice{{\vbox{\hbox{$\textstyle\mkern-1.5mu{-}\mkern-7.5mu{-}$}}\kern-4.675pt}}{{\vbox{\hbox{$\scriptstyle\mkern-1.5mu{-}\mkern-7.5mu{-}$}}\kern-4.02336pt}}{{\vbox{\hbox{$\scriptscriptstyle\mkern-1.5mu{-}\mkern-7.5mu{-}$}}\kern-2.53589pt}}{{\vbox{\hbox{$\scriptscriptstyle\mkern-1.5mu{-}\mkern-7.5mu{-}$}}\kern-1.7709pt}}\!\int\!\!\!\!\!\int_{8Q\cap\Omega}|u-c_{u}|dXdt\leq C\,r\mathchoice{{\vbox{\hbox{$\textstyle\mkern-1.5mu{-}\mkern-7.5mu{-}$}}\kern-4.675pt}}{{\vbox{\hbox{$\scriptstyle\mkern-1.5mu{-}\mkern-7.5mu{-}$}}\kern-4.02336pt}}{{\vbox{\hbox{$\scriptscriptstyle\mkern-1.5mu{-}\mkern-7.5mu{-}$}}\kern-2.53589pt}}{{\vbox{\hbox{$\scriptscriptstyle\mkern-1.5mu{-}\mkern-7.5mu{-}$}}\kern-1.7709pt}}\!\int\!\!\!\!\!\int_{16Q\cap\Omega}|\nabla u|dXdt.$
Since
$\inf_{x\in 2Q\cap\partial\Omega}\widetilde{\mathcal{C}}_{1}(\delta G)(x)\lesssim r^{(-n-1)/p^{\prime}}\|\widetilde{\mathcal{C}}_{1}(\delta G)\|_{L^{p^{\prime}}(\partial\Omega)}\lesssim r^{(-n-1)/p^{\prime}}$
by (4.17), and that $1-\frac{1}{p^{\prime}}=\frac{1}{p}$, we deduce that
$I_{1}\lesssim r^{(n+1)/p}\left(\mathchoice{{\vbox{\hbox{$\textstyle\mkern-1.5mu{-}\mkern-7.5mu{-}$}}\kern-4.675pt}}{{\vbox{\hbox{$\scriptstyle\mkern-1.5mu{-}\mkern-7.5mu{-}$}}\kern-4.02336pt}}{{\vbox{\hbox{$\scriptscriptstyle\mkern-1.5mu{-}\mkern-7.5mu{-}$}}\kern-2.53589pt}}{{\vbox{\hbox{$\scriptscriptstyle\mkern-1.5mu{-}\mkern-7.5mu{-}$}}\kern-1.7709pt}}\!\int\!\!\!\!\!\int_{10Q\cap\Omega}|\nabla u|^{2}\right)^{\frac{1}{2}}$
as desired.
Next, we treat the term $I_{0}$. We define $v$ to be the weak solution to $L^{*}v=-\operatorname{div}G$ in $\Omega$ with zero Neumann data. Then we can write
$I_{0}=\iint_{\Omega}(u-c_{u})\varphi L^{*}v\,dXdt=\iint_{\Omega}(u-c_{u})\varphi\left(\partial_{t}v+\operatorname{div}A^{T}\nabla v\right)\,dXdt\\ =\iint_{\Omega}(u-c_{u})\partial_{t}(\varphi(v-c_{v}))dXdt-\iint_{\Omega}(u-c_{u})(v-c_{v})\,\partial_{t}\varphi\,dXdt-\iint_{\Omega}A\nabla[(u-c_{u})\varphi]\cdot\nabla(v-c_{v})\,dXdt.$
for any constant $c_{v}$ whose value will be determined later. We further write
@@
$I_{0}=\iint_{\Omega}u\partial_{t}(\varphi(v-c_{v}))dXdt-\iint_{\Omega}A\nabla u\cdot\nabla(\varphi(v-c_{v}))dXdt-\iint_{\Omega}(u-c_{u})(v-c_{v})\,\partial_{t}\varphi\,dXdt\\ +\iint_{\Omega}A\nabla u\cdot\nabla\varphi\,(v-c_{v})\,dXdt-\iint_{\Omega}A\nabla\varphi\cdot\nabla v(u-c_{u})dXdt.$
Using the fact that $Lu=0$ in $\Omega\cap KQ$ with zero Neumann data on $\partial\Omega\cap KQ$ and taking $\varphi(v-c_{v})$ as a test function, we obtain that
@@
(4.21) $I_{0}=-\iint_{\Omega}(u-c_{u})(v-c_{v})\,\partial_{t}\varphi\,dXdt+\iint_{\Omega}A\nabla u\cdot\nabla\varphi\,(v-c_{v})\,dXdt-\iint_{\Omega}A\nabla\varphi\cdot\nabla v(u-c_{u})dXdt\\ =:I_{00}+I_{01}+I_{02}.$
To estimate these terms, we first derive the following estimate:
@@
(4.22) $\mathchoice{{\vbox{\hbox{$\textstyle\mkern-1.5mu{-}\mkern-7.5mu{-}$}}\kern-4.675pt}}{{\vbox{\hbox{$\scriptstyle\mkern-1.5mu{-}\mkern-7.5mu{-}$}}\kern-4.02336pt}}{{\vbox{\hbox{$\scriptscriptstyle\mkern-1.5mu{-}\mkern-7.5mu{-}$}}\kern-2.53589pt}}{{\vbox{\hbox{$\scriptscriptstyle\mkern-1.5mu{-}\mkern-7.5mu{-}$}}\kern-1.7709pt}}\!\int\!\!\!\!\!\int_{4Q\cap\Omega}|v-c_{v}|dXdt\lesssim r^{-n-1}\left(\|\widetilde{N}(\delta|\nabla v|{\mathds{1}}_{4KQ})\|_{L^{1}(\partial\Omega)}+\|\widetilde{\mathcal{C}}(\delta G{\mathds{1}}_{4KQ})\|_{L^{1}(\partial\Omega)}\right),$
with a constant $c_{v}$ to be specified shortly.
@@
(4.23) $\iint_{4Q\cap\Omega}|v-c_{v}|dXdt\lesssim\int_{(y,s)\in\partial\Omega}\iint_{\Gamma(y,s)}|v(X,t)-c_{v}|{\mathds{1}}_{4Q}(X,t)\frac{dXdt}{\delta(X,t)^{n+1}}d\sigma(y,s)\\ \lesssim\int_{(y,s)\in\partial\Omega}\sum_{W\in\mathcal{W},\,(y,s)\in 10W}\ell(W)\mathchoice{{\vbox{\hbox{$\textstyle\mkern-1.5mu{-}\mkern-7.5mu{-}$}}\kern-4.675pt}}{{\vbox{\hbox{$\scriptstyle\mkern-1.5mu{-}\mkern-7.5mu{-}$}}\kern-4.02336pt}}{{\vbox{\hbox{$\scriptscriptstyle\mkern-1.5mu{-}\mkern-7.5mu{-}$}}\kern-2.53589pt}}{{\vbox{\hbox{$\scriptscriptstyle\mkern-1.5mu{-}\mkern-7.5mu{-}$}}\kern-1.7709pt}}\!\int\!\!\!\!\!\int_{W}|v(X,t)-c_{v}|{\mathds{1}}_{4Q}(X,t)dXdt\,d\sigma(y,s).$
Let $W_{Q}\in\mathcal{W}$ be a parabolic Whitney cube associated to $Q$, that is, $W_{Q}\subset Q$, $10W_{Q}\cap\partial\Omega\neq\emptyset$, and $\delta(W_{Q},\partial\Omega):=\operatorname{dist}_{p}(W_{Q},\partial\Omega)\gtrsim r$. We take $c_{v}=\mathchoice{{\vbox{\hbox{$\textstyle\mkern-1.5mu{-}\mkern-7.5mu{-}$}}\kern-4.675pt}}{{\vbox{\hbox{$\scriptstyle\mkern-1.5mu{-}\mkern-7.5mu{-}$}}\kern-4.02336pt}}{{\vbox{\hbox{$\scriptscriptstyle\mkern-1.5mu{-}\mkern-7.5mu{-}$}}\kern-2.53589pt}}{{\vbox{\hbox{$\scriptscriptstyle\mkern-1.5mu{-}\mkern-7.5mu{-}$}}\kern-1.7709pt}}\!\int\!\!\!\!\!\int_{W_{Q}}v\,dXdt$.
Fix any $W\in\mathcal{W}$ such that $W\cap 4Q\neq\emptyset$. We construct the Harnack chain of Whitney cubes $W=W_{1},\dots,W_{M}=W_{Q}$. Such Harnack chain satisfies
@@
    for any $1\leq i\leq M$ we have $2W_{i}\in\gamma^{*}(y,s)$ for $(y,s)\in 10W\cap\partial\Omega$ and a parabolic cone $\gamma^{*}(y,s)$ with large enough aperture since $(y,s)\in 10W$;

### v2

@@ 3. Basic estimates
The proof follows the strategy of [13, Lemma 4.1], adapted to the parabolic setting. Fix $J_{r}:=J_{r}(\bar{x},\bar{t})$ with $(\bar{x},\bar{t})\in Q_{0}\cap\partial\Omega$ and $J_{2r}:=J_{2r}(\bar{x},\bar{t})\subset Q_{0}$. We claim that we can assume that for some $K\geq 1$ (independent of $J_{r}$ and $u$) to be determined during the proof, $J_{Kr}(\bar{x},\bar{t})\subset Q_{0}$, and that the desired estimate (4.16) would follow once we prove that
(4.17) $\|\widetilde{N}(|\nabla u|{\mathds{1}}_{J_{r}})\|_{L^{p}(\partial\Omega)}\lesssim r^{(n+1)/p}\Big(\mathchoice{{\vbox{\hbox{$\textstyle\mkern-1.5mu{-}\mkern-7.5mu{-}$}}\kern-4.675pt}}{{\vbox{\hbox{$\scriptstyle\mkern-1.5mu{-}\mkern-7.5mu{-}$}}\kern-4.02336pt}}{{\vbox{\hbox{$\scriptscriptstyle\mkern-1.5mu{-}\mkern-7.5mu{-}$}}\kern-2.53589pt}}{{\vbox{\hbox{$\scriptscriptstyle\mkern-1.5mu{-}\mkern-7.5mu{-}$}}\kern-1.7709pt}}\!\int\!\!\!\!\!\int_{J_{Kr}\cap\Omega}|\nabla u|^{2}dXdt\Big)^{\frac{1}{2}}.$
In fact, if (4.17) holds, then we cover $J_{r}$ by a finite collection of backward parabolic cubes $J_{i}:=J_{r_{i}}(x_{i},t_{i})$ with $r_{i}\approx r$ such that either
(i) $(x_{i},t_{i})\in\partial\Omega$ and $J_{Kr_{i}}(x_{i},t_{i})\subset J_{2r}$, or
(ii) $J_{2r_{i}}(x_{i},t_{i})\subset J_{2r}\cap\Omega$.
As there are finitely many $J_{i}$ and the number depends only on the dimension, we have that
$\|\widetilde{N}(|\nabla u|{\mathds{1}}_{J_{r}})\|_{L^{p}(\partial\Omega)}\lesssim\sup_{i}\|\widetilde{N}(|\nabla u|{\mathds{1}}_{J_{i}})\|_{L^{p}(\partial\Omega)}.$
If $J_{i}$ is as in (i), then we use (4.17) to get that
$\|\widetilde{N}(|\nabla u|{\mathds{1}}_{J_{i}})\|_{p}\lesssim r_{i}^{(n+1)/p}\left(\mathchoice{{\vbox{\hbox{$\textstyle\mkern-1.5mu{-}\mkern-7.5mu{-}$}}\kern-4.675pt}}{{\vbox{\hbox{$\scriptstyle\mkern-1.5mu{-}\mkern-7.5mu{-}$}}\kern-4.02336pt}}{{\vbox{\hbox{$\scriptscriptstyle\mkern-1.5mu{-}\mkern-7.5mu{-}$}}\kern-2.53589pt}}{{\vbox{\hbox{$\scriptscriptstyle\mkern-1.5mu{-}\mkern-7.5mu{-}$}}\kern-1.7709pt}}\!\int\!\!\!\!\!\int_{J_{Kr_{i}}(x_{i},r_{i})\cap\Omega}\left|\nabla u\right|^{2}\right)^{1/2}\lesssim r^{(n+1)/p}\left(\mathchoice{{\vbox{\hbox{$\textstyle\mkern-1.5mu{-}\mkern-7.5mu{-}$}}\kern-4.675pt}}{{\vbox{\hbox{$\scriptstyle\mkern-1.5mu{-}\mkern-7.5mu{-}$}}\kern-4.02336pt}}{{\vbox{\hbox{$\scriptscriptstyle\mkern-1.5mu{-}\mkern-7.5mu{-}$}}\kern-2.53589pt}}{{\vbox{\hbox{$\scriptscriptstyle\mkern-1.5mu{-}\mkern-7.5mu{-}$}}\kern-1.7709pt}}\!\int\!\!\!\!\!\int_{J_{2r}\cap\Omega}\left|\nabla u\right|^{2}\right)^{1/2}.$
If $J_{i}$ is as in (ii), then we observe that $\widetilde{N}(|\nabla u|{\mathds{1}}_{J_{i}})$ is supported on a large shadow of $J_{i}$ on $\partial\Omega$, which is of measure $Cr^{n+1}$. Therefore,
$\|\widetilde{N}(|\nabla u|{\mathds{1}}_{J_{i}})\|_{L^{p}(\partial\Omega)}\approx r^{(n+1)/p}\left(\mathchoice{{\vbox{\hbox{$\textstyle\mkern-1.5mu{-}\mkern-7.5mu{-}$}}\kern-4.675pt}}{{\vbox{\hbox{$\scriptstyle\mkern-1.5mu{-}\mkern-7.5mu{-}$}}\kern-4.02336pt}}{{\vbox{\hbox{$\scriptscriptstyle\mkern-1.5mu{-}\mkern-7.5mu{-}$}}\kern-2.53589pt}}{{\vbox{\hbox{$\scriptscriptstyle\mkern-1.5mu{-}\mkern-7.5mu{-}$}}\kern-1.7709pt}}\!\int\!\!\!\!\!\int_{J_{i}}|\nabla u|^{2}\,dX\right)^{\frac{1}{2}}\approx r^{(n+1)/p}\left(\mathchoice{{\vbox{\hbox{$\textstyle\mkern-1.5mu{-}\mkern-7.5mu{-}$}}\kern-4.675pt}}{{\vbox{\hbox{$\scriptstyle\mkern-1.5mu{-}\mkern-7.5mu{-}$}}\kern-4.02336pt}}{{\vbox{\hbox{$\scriptscriptstyle\mkern-1.5mu{-}\mkern-7.5mu{-}$}}\kern-2.53589pt}}{{\vbox{\hbox{$\scriptscriptstyle\mkern-1.5mu{-}\mkern-7.5mu{-}$}}\kern-1.7709pt}}\!\int\!\!\!\!\!\int_{J_{2r}\cap\Omega}|\nabla u|^{2}\right)^{\frac{1}{2}}.$
Combining both cases, we have shown
$\|\widetilde{N}(|\nabla u|{\mathds{1}}_{J_{r}})\|_{L^{p}(\partial\Omega)}\leq Cr^{(n+1)/p}\left(\mathchoice{{\vbox{\hbox{$\textstyle\mkern-1.5mu{-}\mkern-7.5mu{-}$}}\kern-4.675pt}}{{\vbox{\hbox{$\scriptstyle\mkern-1.5mu{-}\mkern-7.5mu{-}$}}\kern-4.02336pt}}{{\vbox{\hbox{$\scriptscriptstyle\mkern-1.5mu{-}\mkern-7.5mu{-}$}}\kern-2.53589pt}}{{\vbox{\hbox{$\scriptscriptstyle\mkern-1.5mu{-}\mkern-7.5mu{-}$}}\kern-1.7709pt}}\!\int\!\!\!\!\!\int_{J_{2r}\cap\Omega}|\nabla u|^{2}dXdt\right)^{\frac{1}{2}},$
as desired.
We choose a piecewise smooth function $\varphi$ so that $\varphi\equiv 1$ on $J_{\frac{3}{2}r}$, $\varphi\equiv 0$ outside $J_{2r}$, $0\leq\varphi\leq 1$, $|\nabla\varphi|\lesssim\frac{1}{r}$ and that $|\partial_{t}\varphi|\lesssim\frac{1}{r^{2}}$ in $J_{2r}$. By duality, i.e. (2.15), there exists $G\in L^{\infty}_{c}(\Omega,\mathbb{R}^{n})$ such that
@@
(4.19) $\|\widetilde{N}(\nabla u{\mathds{1}}_{J_{r}})\|_{L^{p}(\partial\Omega)}\leq C\iint_{\Omega}{\mathds{1}}_{J_{r}}\nabla u\cdot G\,dXdt$
Let $c_{u}$ be a constant to be determined later. We have that
@@
$\|\widetilde{N}(\nabla u{\mathds{1}}_{J_{r}})\|_{L^{p}(\partial\Omega)}\lesssim\iint_{\Omega}{\mathds{1}}_{J_{r}}\nabla u\cdot G\,\varphi\,dXdt=\iint_{\Omega}{\mathds{1}}_{J_{r}}\nabla(u-c_{u})\cdot G\,\varphi\,dXdt\\ =\int_{T_{0}}^{\bar{t}}\int_{\mathcal{O}}{\mathds{1}}_{J_{r}}\nabla\left((u-c_{u})\varphi\right)\cdot G\,dXdt-\iint_{J_{r}\cap\Omega}(u-c_{u})\nabla\varphi\cdot G\,dXdt\\ =\int_{T_{0}}^{\bar{t}}\int_{\mathcal{O}}{\mathds{1}}_{J_{r}}\nabla\left((u-c_{u})\varphi\right)\cdot G\,dXdt=:I_{0},$
@@
where we have used the observation that $\iint_{J_{r}\cap\Omega}(u-c_{u})\nabla\varphi\cdot G\,dXdt=0$ since $\nabla\varphi=0$ in $J_{\frac{3}{2}r}\supset J_{r}$.
To treat the term $I_{0}$, we define $v\in\dot{E}$ to be the energy solution to $L^{*}v=-\operatorname{div}(G{\mathds{1}}_{J_{r}})$ in $\Omega$ with zero Neumann data. We take $(u-c_{u})\varphi$ as a test function in the PDE satisfied by vv—which is valid since $(u-c_{u})\varphi\in L^{2}(\mathbb{R},\mathcal{W}^{1,2}(Q_{0}^{\prime}\cap\mathcal{O}))$; see also (2.27) and the remark below—and integrate in time from $T_{0}$ to $\bar{t}$ to get that [^4] We shall be somewhat imprecise regarding the domain of integration in the spatial variables: we write the integral over $\mathcal{O}$, although $u$ is only defined on $Q_{0}\cap\mathcal{O}$. This causes no problem, since we may extend $u$ arbitrarily outside $Q_{0}$, as the cutoff function $\varphi(\cdot,t)$ is compactly supported in $J_{2r}\cap\mathcal{O}\subset Q_{0}\cap\mathcal{O}$.
@@
$I_{0}=\int_{T_{0}}^{\bar{t}}\langle\partial_{t}v(\cdot,t),(u(\cdot,t)-c_{u})\varphi(\cdot,t)\rangle_{W^{-1,2},\mathcal{W}^{1,2}}dt-\int_{T_{0}}^{\bar{t}}\int_{\mathcal{O}}A^{T}\nabla v\cdot\nabla[(u-c_{u})\varphi]dXdt\\ =\int_{T_{0}}^{\bar{t}}\langle\partial_{t}v(\cdot,t),(u(\cdot,t)-c_{u})\varphi(\cdot,t)\rangle_{W^{-1,2},\mathcal{W}^{1,2}}dt-\int_{T_{0}}^{\bar{t}}\int_{\mathcal{O}}A\nabla[(u-c_{u})\varphi]\cdot\nabla v\,dXdt.$
We want to use the PDE satisfied by $u-c_{u}$, and so we further rewrite the second term using the product rule:
$\int_{T_{0}}^{\bar{t}}\int_{\mathcal{O}}A\nabla[(u-c_{u})\varphi]\cdot\nabla v\,dXdt=-\int_{T_{0}}^{\bar{t}}\int_{\mathcal{O}}A\nabla(u-c_{u})\cdot\nabla(\varphi\,v)dXdt\\ +\int_{T_{0}}^{\bar{t}}\int_{\mathcal{O}}A\nabla(u-c_{u})\cdot\nabla\varphi\,v\,dXdt-\int_{T_{0}}^{\bar{t}}\int_{\mathcal{O}}A\nabla\varphi\cdot\nabla v\,(u-c_{u})dXdt.$
For the first term coming from $I_{0}$, we claim that
(4.20) $\int_{T_{0}}^{\bar{t}}\langle\partial_{t}v(\cdot,t),(u(\cdot,t)-c_{u})\varphi(\cdot,t)\rangle_{\mathcal{W}^{-1,2},\mathcal{W}^{1,2}}dt\\ =-\int_{T_{0}}^{\bar{t}}\langle\partial_{t}(u-c_{u}),v(\cdot,t)\varphi(\cdot,t)\rangle_{\mathcal{W}^{-1,2},\mathcal{W}^{1,2}}dt-\int_{T_{0}}^{\bar{t}}\int_{\mathcal{O}}(u-c_{u})v\,\partial_{t}\varphi\,dXdt.$
In fact, by a density argument, one can assume that $u,v\in C^{\infty}(\mathcal{O}\times(T_{0},T_{1}))$ and then the pairing of $\mathcal{W}^{-1,2}(\mathcal{O})$ and $\mathcal{W}^{1,2}(\mathcal{O})$ becomes inner product in $L^{2}(\mathcal{O})$. In addition, observe that since $v$ is a solution to $L^{*}v=0$ in $\mathcal{O}\times[\bar{t},\infty)$ with 0 Neumann data, $v$ is constant in $\left\{t\geq\bar{t}\right\}$ (notice that we use here that $L^{*}$ is backward in time parabolic PDE). Without loss of generality, we can assume that the constant is 0, that is,
(4.21) $v(X,t)=0\qquad\text{for $t\geq\bar{t}$}.$
Therefore, integrating by parts in $t$, the left-hand side of (4.20) can be written as
$\int_{T_{0}}^{\bar{t}}\int_{\mathcal{O}}\partial_{t}v(u-c_{u})\varphi\,dXdt\\ =\int_{\mathcal{O}}\int_{T_{0}}^{\bar{t}}\partial_{t}\left(v(u-c_{u})\varphi\right)dtdX-\int_{\mathcal{O}}\int_{T_{0}}^{\bar{t}}\partial_{t}(u-c_{u})v\,\varphi\,dtdX-\int_{T_{0}}^{\bar{t}}\int_{\mathcal{O}}(u-c_{u})v\,\partial_{t}\varphi\,dXdt\\ =-\int_{T_{0}}^{\bar{t}}\int_{\mathcal{O}}\partial_{t}(u-c_{u})v\,\varphi\,dtdX-\int_{T_{0}}^{\bar{t}}\int_{\mathcal{O}}(u-c_{u})v\,\partial_{t}\varphi\,dXdt$
thanks to (4.21).
Since $u-c_{u}$ satisfies $Lu=0$ in $Q_{0}\cap\Omega$ with zero Neumann data on $\partial\Omega\cap Q_{0}$ (recall that $J_{Kr}\subset Q_{0}$), we take $\varphi\,v$ as a test function and integrate in time from $T_{0}$ to $\bar{t}$ to get that
$\int_{T_{0}}^{\bar{t}}\langle\partial_{t}(u-c_{u}),v(\cdot,t)\varphi(\cdot,t))\rangle_{\mathcal{W}^{-1,2},\mathcal{W}^{1,2}}dt+\int_{T_{0}}^{\bar{t}}\int_{\mathcal{O}}A\nabla(u-c_{u})\cdot\nabla(\varphi\,v)dXdt=0.$
Therefore,
(4.22) $I_{0}=-\int_{T_{0}}^{\bar{t}}\int_{\mathcal{O}}(u-c_{u})v\,\partial_{t}\varphi\,dXdt+\int_{T_{0}}^{\bar{t}}\int_{\mathcal{O}}A\nabla u\cdot\nabla\varphi\,v\,dXdt\\ -\int_{T_{0}}^{\bar{t}}\int_{\mathcal{O}}A\nabla\varphi\cdot\nabla v(u-c_{u})dXdt=:I_{00}+I_{01}+I_{02}.$
To estimate these terms, we first derive the following estimate: let $Q:=Q_{r}(\bar{x},\bar{t})$ and $kQ:=Q_{kr}(\bar{x},\bar{t})$ for $k>0$,
@@
(4.23) $\mathchoice{{\vbox{\hbox{$\textstyle\mkern-1.5mu{-}\mkern-7.5mu{-}$}}\kern-4.675pt}}{{\vbox{\hbox{$\scriptstyle\mkern-1.5mu{-}\mkern-7.5mu{-}$}}\kern-4.02336pt}}{{\vbox{\hbox{$\scriptscriptstyle\mkern-1.5mu{-}\mkern-7.5mu{-}$}}\kern-2.53589pt}}{{\vbox{\hbox{$\scriptscriptstyle\mkern-1.5mu{-}\mkern-7.5mu{-}$}}\kern-1.7709pt}}\!\int\!\!\!\!\!\int_{4Q\cap\Omega}|v|dXdt\lesssim r^{-n-1}\left(\|\widetilde{N}(\delta|\nabla v|{\mathds{1}}_{4KQ})\|_{L^{1}(\partial\Omega)}+\|\widetilde{\mathcal{C}}(\delta G{\mathds{1}}_{4KQ})\|_{L^{1}(\partial\Omega)}\right).$
@@
(4.24) $\iint_{4Q\cap\Omega}|v|dXdt\lesssim\int_{(y,s)\in\partial\Omega}\iint_{\Gamma(y,s)}|v(X,t)|{\mathds{1}}_{4Q}(X,t)\frac{dXdt}{\delta(X,t)^{n+1}}d\sigma(y,s)\\ \lesssim\int_{(y,s)\in\partial\Omega}\sum_{W\in\mathcal{W},\,(y,s)\in 10W}\ell(W)\mathchoice{{\vbox{\hbox{$\textstyle\mkern-1.5mu{-}\mkern-7.5mu{-}$}}\kern-4.675pt}}{{\vbox{\hbox{$\scriptstyle\mkern-1.5mu{-}\mkern-7.5mu{-}$}}\kern-4.02336pt}}{{\vbox{\hbox{$\scriptscriptstyle\mkern-1.5mu{-}\mkern-7.5mu{-}$}}\kern-2.53589pt}}{{\vbox{\hbox{$\scriptscriptstyle\mkern-1.5mu{-}\mkern-7.5mu{-}$}}\kern-1.7709pt}}\!\int\!\!\!\!\!\int_{W}|v(X,t)|{\mathds{1}}_{4Q}(X,t)dXdt\,d\sigma(y,s).$
Let $W_{Q}\in\mathcal{W}$ be a parabolic Whitney cube associated to $Q$, that is, $W_{Q}\subset Q$, $10W_{Q}\cap\partial\Omega\neq\emptyset$, and $\delta(W_{Q},\partial\Omega):=\operatorname{dist}_{p}(W_{Q},\partial\Omega)\gtrsim r$. Moreover, we choose $W_{Q}\subset\left\{t>\bar{t}\right\}$, and so $\mathchoice{{\vbox{\hbox{$\textstyle\mkern-1.5mu{-}\mkern-7.5mu{-}$}}\kern-4.675pt}}{{\vbox{\hbox{$\scriptstyle\mkern-1.5mu{-}\mkern-7.5mu{-}$}}\kern-4.02336pt}}{{\vbox{\hbox{$\scriptscriptstyle\mkern-1.5mu{-}\mkern-7.5mu{-}$}}\kern-2.53589pt}}{{\vbox{\hbox{$\scriptscriptstyle\mkern-1.5mu{-}\mkern-7.5mu{-}$}}\kern-1.7709pt}}\!\int\!\!\!\!\!\int_{W_{Q}}v\,dXdt=0$ in view of (4.21).
Fix any $W\in\mathcal{W}$ such that $W\cap 4Q\neq\emptyset$. We construct a Harnack chain of Whitney cubes $W=W_{1},\dots,W_{M}=W_{Q}$. Such a Harnack chain satisfies
@@
    for any $1\leq i\leq M$ we have $2W_{i}\in\gamma^{*}(y,s)$ for any $(y,s)\in 10W\cap\partial\Omega$ and any parabolic cone $\gamma^{*}(y,s)$ with large enough aperture;